OpenStudy (b77w):

Complex number in standard form. How do I do this? 5/(1+i)

OpenStudy (amistre64):

is a+bi standard form? if so you might want to run a conjugate

OpenStudy (b77w):

Yes

OpenStudy (amistre64):

then a conjugate it is: recall that (a+b)(a-b)=a^2-b^2

OpenStudy (b77w):

ok now what do i do?

OpenStudy (amistre64):

essentially you mulitply by 1, but in the form of the conjugate ....

OpenStudy (amistre64):

\[\frac{k}{a+b}*1\] \[\frac{k}{a+b}*\frac{a-b}{a-b}\] \[\frac{ka-kb}{a^2-b^2}\] \[\frac{ka}{a^2-b^2}-\frac{kb}{a^2-b^2}\]

OpenStudy (b77w):

What is k?

OpenStudy (anonymous):

It's just a general variable. Here's what you'll be doing. Take your denominator and flip the sign between the 1 and the i. This is your conjugate, in this case it's (1-i). Now multiply the numerator and denominator by (1+i)

OpenStudy (b77w):

(5-5i)/2

OpenStudy (b77w):

Right?

OpenStudy (anonymous):

Right, exactly, so thats your answer: 2.5 - 2.5i

OpenStudy (b77w):

Thank you so much!

OpenStudy (anonymous):

No problem, glad I could help